OpenStudy (anonymous):

16^(x+4)=8^(3x-2) Solve

OpenStudy (kirbykirby):

Convert it into base 2 first: \[\large 2^{4(x+4)}=2^{3(3x-2)}\\ \large 2^{4x+16}=2^{9x-6}\] Log both sides: \[\large \log \left(2^{4x+16}\right)=\log\left(2^{9x-6}\right)\\ \large (4x+16) \log 2 = (9x-6) \log 2\\ \large (4x)\log 2 + 16 \log 2 = (9x)\log 2 - 6 \log 2\\ \large (4x) \log 2 - (9x) \log 2 = -16 \log 2 - 6 \log 2 \\ \large x\left(4 \log 2 - 9 \log 2 \right) = - 22 \log 2 \\ \large x \left( -5 \log 2 \right) = - 22 \log 2 \\ \large x = \frac{-22 \log 2 }{- 5 \log 2}\\ \large x = \frac{22}{5}\]